### v2

# Career Modeling with Missing Data and Traces

Abstract
Many social scientists study the career trajectories of populations of interest, such as economic and administrative elites. However, data to document such processes are rarely completely available, which motivates the adoption of inference tools that can account for large numbers of missing values. Taking the example of public-private paths of elite civil servants in France, we introduce binary Markov switching models to perform Bayesian data augmentation. Our procedure relies on two data sources: (1) detailed observations of a small number of individual trajectories, and (2) less informative “traces” left by all individuals, which we model for imputation of missing data. An advantage of this model class is that it maintains the properties of hidden Markov models and enables a tailored sampler to target the posterior, while allowing for varying parameters across individuals and time. We provide two applied studies which demonstrate this can be used to properly test substantive hypotheses, and expand the social scientific literature in various ways. We notably show that the rate at which ENA graduates exit the French public sector has not increased since 1990, but that the rate at which they come back has increased.
Keywords: Bayesian Markov Switching Model, Career Modeling, Public-Private Paths, Civil Servants, Missing Values
1 Introduction
Many important questions in the social sciences require understanding career trajectories of certain populations. This is an important aspect when describing political (Jäckle and Kerby, 2018), economic (Dudouet et al., 2010), academic (Angermuller and Blanchard, 2023) and administrative (Rouban, 2014) elites, but also when studying specific professions like musicians (Abbott and Hrycak, 1990), journalists (Machut and Bastin, 2023), or even athletes (Li et al., 2018). Studying careers is seen as interesting in its own right, but also to describe broader social mechanisms such as lobbying and revolving doors (Blanes i Vidal et al., 2012) or the spatial structuring of finance (Bühlmann et al., 2024).
However, a key problem in career modeling is data missingness. It is very rare that all data relevant to the career of a certain population are available in a public record or can even be collected because of resource constraints, which is problematic for two reasons. First, for a given defined population, inference can be made more difficult and less powerful if a majority of individuals cannot be observed. This reduces the breadth of questions that can be answered using career modeling, as it makes the testing of interesting interaction effects impossible in practice. Second, researchers often choose to only study populations that can be entirely observed. This forces them to study the groups that are the most visible, such as top political elites; it thus constrains the breadth of questions that can be addressed using career modeling.
Our article has two aims. First, we propose a general methodology employing hidden Markov models with covariates and autoregressive components, and showcase, using plausible synthetic data examples, why they are required for such a modeling task while remaining simple enough for substantive hypothesis testing. The procedure relies on (i) a fraction of the population being well-observed using a first (possibly hand collected) data source, and (ii) the usage of partial “traces” from a digital or administrative record that are modeled to input the missing information for the remaining majority of individuals. Second, we provide two applied studies showcasing the usefulness of these models, which expand the social scientific literature of administrative elites in meaningful ways.
Our leading example will be the study of public-private paths of elite civil servants in France. Paths of civil servants into the private sector are seen as important both by the general public and in the social science literature, where many publications that quantitatively study French high civil servants tackle the issue directly (Bouzidi et al., 2010; Kolopp, 2021; Rouban, 2014). Public-private paths broadly understood for policy-relevant positions are sometimes described using the term “revolving doors”, because these back-and-forth switches resemble a physical revolving door (Seabrooke and Tsingou, 2021; France and Vauchez, 2017), but other scholars define revolving doors to be a much more precise and limited in scope phenomenon. Unassuming about this definitional debate, we denote this problem as the study of public-private paths of French elite civil servants, and acknowledge that our article falls in the first category of works.
The phenomenon of public-private paths of civil servants demonstrates well the usefulness of the general method because it shares many difficulties with other settings in career analysis. First, it is relatively easy to describe the phenomenon for a portion of elite civil servants, by collecting data from digital networking websites such as LinkedIn on which they are registered. However, a majority of them do not have a networking digital profile, and further, it is not always the case that a profile is fully completed from the starting point of one’s career. Second, most civil servants, to various degrees, leave “traces” of their activity, when they sign legal documents, or obtain important promotions or bonuses. We showcase that it is possible to reconstruct trajectories of large civil servant populations and draw interpretations that contribute to the social scientific literature. In addition, our population is voluntarily more diverse and an order of magnitude larger compared to usual prosopography studies of the domain, to demonstrate that our methodology is well suited for analyses that are also targeted at less visible, larger professional groups.
We introduce and discuss in Section 2 the general statistical framework for career modeling with partially observed data, then apply it on real data to conduct two studies on French elite civil servants in Section 3. Section 4 concludes both on the general methodology and on the substantive learning exercise on civil servants.
2 General methodology
2.1 Problem setting
2.1.1 Detailed and trace data sources
The core of our approach stems from the tension between data quality and quantity. In the majority of cases, it is difficult to obtain data that exhibit both high precision to document career trajectories, and high coverage. However, there often exists at least two data sources with opposed strengths and weaknesses:
  • Detailed data source. One data source exhibits low coverage but is very informative on the individuals it documents. This usually corresponds to data collected by hand for a small number of individuals, but can also be taken automatically from digital or administrative sources, such as professional networking websites.
  • Trace data source. Another data source exhibits high coverage but is not very informative taken on its own and requires statistical elaboration for analysis. This concerns larger administrative or digital records, which we will describe as collections of “traces”: we know that individuals in a certain career position should regularly leave traces of their activity in these records, but the rate at which this happens varies across people and time.
Let us provide several examples taken from the introduction. When studying journalists’ affiliations, we know that some of them have digital profiles on networking websites (detailed data source), or we can spend some time hand-crafting a career database on a subpopulation; but we also know that journalists regularly publish newspaper articles with their name and affiliation (trace data source). We can apply the same logic to academic researchers, even if the rate at which they publish is lower: modeling the process by which individuals emit traces needs to be informed by data, and this motivates our mathematical framework. For board members in the economic elite, one might hand-construct a dataset (detailed data source); but we might also find their names with affiliation in documents at the end of every board meeting (trace data source), which might vary in frequency from company to period. Our leading example will be the case of elite civil servants in the public and private sectors: a good portion of them have a digital presence (detailed data source); but for many groups, we know they leave traces of their activity in administrative records when they get certain promotions or take certain decisions in the public sector (trace data source). For all the examples cited in the introduction, one could find credible data sources that would match this setting.
2.1.2 Formalization
We now turn to the formalization of this setting. We aim to study the career trajectory of $N$ individuals $i\in\{1,...,N\}$ in discrete time $t\in\{1,...,T\}$. We denote the career position of individual $i$ in time $t$ as $X_{i,t}\in\mathcal{X}$, with $\mathcal{X}=\{x_{1},...,x_{K}\}$ a known finite state space of size $K$. In our application, $\mathcal{X}$ is equal to $\{0,1\}$, which corresponds to public and private sectors, but this can describe many other settings: position in a given newsroom or media type for a journalist, stage of the career for an academic, company for a corporate manager… As we will use a model in the family of Hidden Markov Models (HMMs), we expect that choosing a higher cardinality for $\mathcal{X}$ will incur further assumptions so that the parameters governing the behavior remain of moderate effective dimension. The ensemble of all trajectories at all time points is denoted $\textbf{X}=(X_{i,t})_{i,t}$. Observe that time here is discrete: our techniques may be generalized with continuous time, but we do not find a sufficient interest given the prohibitive cost that would incur. In our applications, $t$ corresponds to a six-month period; other choices could be reasonable depending on the substantive research question at hand. We assume that X is partially observed, thanks to the detailed data source: for some (but not all) values of $i$ and $t$, we observe the value of $X_{i,t}$.
We now need to connect it to information found in the trace data source. The trace data source is described using a second set of time series, $\textbf{Y}=(Y_{i,t})_{i,t}$, which take value $Y_{i,t}\in\mathcal{Y}$ for each $i$ at time $t$. The state space $\mathcal{Y}$ may here be either discrete or continuous, one-dimensional or multi-dimensional. In our application, it will simply be a binary variable describing whether an administrative trace can be found for individual $i$ at time $t$. This could correspond to the number of articles published by a journalist or an academic, an encoding of an athlete’s participation in different competitions, or a binary encoding on whether there was a document including the name of a board member in the given period. Unlike X, we assume that Y is fully observed.
Lastly, this setting is described with covariates $\textbf{W}=(W_{i,t})_{i,t}$ which are of interest to the researchers or which are deemed important to control for, and which may vary across individuals and time. A typical example will be to include the sex of individuals, their age, or their initial diplomas, to see if career behavior differs across these characterics. Importantly, these covariates will be used both to model the behavior of career paths X, and the relationship between paths X and traces Y, and so may include both substantive and nuisance characteristics. We will show it is possible and important to include transformations of lagged values of $Y_{i,t}$ into $W_{i,t}$ when describing the relationship between X and Y.
2.2 Model building with Markov switching models and inference
2.2.1 Related works and inspirations
Our modeling approach is inspired by duration analysis models and their connection to sequence analysis, and by recent developments using Markov switching models in realistic applied settings.
When working with social sequence data, many questions are fundamentally tied to duration. For example, duration analysis models have been applied the durability of political elites such as ministers (King et al., 1990; Bienen and Van de Walle, 1991). In addition, many applications find interest in the duration individuals may stay in more than one type of position: Machut and Bastin (2023) are interested in the career instability of journalists and how transitions across steps of the careers are accelerating with time; Angermuller and Blanchard (2023) are interested in the mobility of academic researchers across statuses, institutions, and countries; and we are interested in both the time elite civil servants spend in the public sector and the time they take before coming back. In our specific application, we complement the scholarship on the tenure of administrative elites using traditional duration analysis (Fleischer, 2016) since we also consider the time taken before coming back, and similar generalizations for the other examples can be provided by modeling as a whole trajectories X instead of only focusing on a single duration problem.
One traditional solution is to study discrete trajectories as Markov models, or, in the case there are missing values, as hidden Markov models (HMM). A complication is that we aim to retain a traditional strength of duration analyses, namely, the ability to flexibly include covariates whose importance can be assessed by statistical tests, as in proportional hazard models (Box-Steffensmeier et al., 2007). A second inspiration of this work thus consists of using hidden Markov models with covariates, meaning that the transition rates of X and emission probabilities of Y can vary across individuals and time. This class of models, otherwise called Markov switching, has been regularly used in recent applied statistics works. Wang and Puterman (1999) used Markov Poisson regression models for discrete time series, and Langrock et al. (2017) proposed generalized additive models with B-splines for additional covariate flexibility. Recent applications include financial and economic forecasting (Meligkotsidou and Dellaportas, 2011), biometrics and environmental data (Wong, 2001), and shooting performance in sports (Sandri et al., 2020). These applications exhibit how Markov switching models produce good estimates yet remain easy to handle.
2.2.2 Formal presentation
For each individual $i$, we assume that the trajectory $(X_{t}^{i})_{t}$ is a Markov chain with non-homogeneous transition kernel $P_{i,t}=(\gamma_{i,t}^{x,x^{\prime}})_{x,x^{\prime}\in\mathcal{X}}$. For each starting point $x$, we model the transition probabilities using a generalized linear model (GLM) for the multinomial distribution:
$(\gamma_{i,t}^{x,x_{1}},...,\gamma_{i,t}^{x,x_{K}})=\text{softmax}((W_{i,t}^{T}\beta^{x,x_{1}},...,W_{i,t}^{T}\beta^{x,x_{K}})),$ (1)
with $W_{i,t}$ the aforementioned vector of covariates, and $(\beta^{x,x^{\prime}})_{x^{\prime}\in\mathcal{X}}$ vectors of parameters. Notice that we may have different parameters $\beta^{x,x^{\prime}}$ for different transitions $(x,x^{\prime})\in\mathcal{X}\times\mathcal{X}$, but that the parameters $\beta^{x,x^{\prime}}$ do not depend on time. Time is rather encoded here as a variable in the covariate vectors $W_{i,t}$. Notice also that, for a state space of size $K$, there are only $K\times(K-1)$ vectors of parameters $(\beta^{x,x^{\prime}})$ to model. In the specific case of career modeling, where individuals do not move at every step, it is reasonable to take the absence of change $x=x^{\prime}$ as the reference category.
For every point $x\in\mathcal{X}$, we next need to model the trace-emission process, that is the distribution of $Y_{i,t}$ conditional on $X_{i,t}=x$. We will use a parametric family with likelihood $\mathcal{H}$ and parameter $\lambda_{i,t}$ specified via a GLM model:
$\lambda_{i,t}:=\mathbb{E}[Y_{i,t}|X_{i,t}=x,\widetilde{W}_{i,t}]=g(\widetilde{W}_{i,t}^{T}\eta^{x}),$ (2)
with $\widetilde{W}_{i,t}$ the vector of covariates, $(\eta^{x})_{x\in\mathcal{X}}$ vectors of parameters, and $g$ a link function. We stress that we use the notation $\widetilde{W}_{i,t}$ to communicate that when estimating emission probabilities, it is possible to include lagged values of $(Y_{i,t^{\prime}})_{t^{\prime}<t}$ inside $\widetilde{W}_{i,t}$, in addition to the standard covariates in ${W}_{i,t}$. Formally, $\widetilde{W}_{i,t}=[W_{i,t}^{T},f(Y_{i,1},...,Y_{i,t-1})]^{T}$, with $f$ any user-specified transformation. We discuss why this auto-correlation feature is essential for such applications in subsection 2.2.3.
Figure 1: Comparison between standard HMM and Markov switching model.
A graphical presentation of these assumptions and a comparison with traditional HMM are available in Figure 1. The likelihood is available in closed form for all parameter vectors for transitions $(\beta^{x,x^{\prime}})_{x,x^{\prime}\in\mathcal{X}}$:
$\forall x\in\mathcal{X},\mathcal{L}(\beta^{x,.}|\textbf{X, W})=\prod_{i=1}^{N}\prod_{\begin{subarray}{c}t=1\\ X_{i,t}=x\end{subarray}}^{T-1}\prod_{x^{\prime}\in\mathcal{X}}\left(\gamma_{i,t}^{x,x^{\prime}}\right)^{1_{\{X_{i,t+1}=x^{\prime}\}}}$ (3)
and for emissions $(\eta^{x})_{x\in\mathcal{X}}$
$\forall x\in\mathcal{X},\mathcal{L}(\eta^{x}|\textbf{X, Y, W})=\prod_{i=1}^{N}\prod_{\begin{subarray}{c}t=1\\ X_{i,t}=x\end{subarray}}^{T}\mathcal{H}(Y_{i,t}|\lambda_{i,t}^{x}),$ (4)
where $\mathcal{H}$ is the likelihood function for the trace-emission GLM model, which will be practical for exploring the space of parameters with a custom sampler.
We purposefully introduced the Markov switching model in a general fashion, but it is now useful to provide details on one useful specificity of applications for career modeling. In many instances, certain values taken by $X_{i,t}\in\mathcal{X}$ will be tied to certain values or a certain support of $Y_{i,t}$, deterministically. For example: when studying traces in the public sector record, we know that if an individual is working in the private sector (denoted here as $X_{i,t}=\text{Private}$) then there cannot be any traces in the public sector record. Formally, $X_{i,t}=\text{Private}$ implies $Y_{i,t}=0$. This can be applied to many cases where individuals cannot hold multiple positions at the same time: journalists who have a full-time contract, managers in companies, politicians and their political party.
There is no need to actively model the emission parameters $\eta^{x}$ in cases $x\in\mathcal{X}$ where we have such information (e.g. $\eta^{Private}$ for the civil servants), since, because of our external knowledge, the conditional distribution of $Y_{i,t}$ given $X_{i,t}=x$ is deterministic. This reduces the number of parameters to estimate and enables the seamless integration of expert knowledge. Additionally, when trying to reconstruct trajectories X from traces Y, this will provide an unusually strong conditioning compared to usual applications of HMMs. This is a reason why our procedure works well in practice. As a whole, this application specificity thus facilitates both inference of parameters and reconstruction of trajectories.
2.2.3 Motivation from a complexity point of view
The model we suggest is purposefully more complex than a Hidden Markov Model (HMM). Indeed, several issues arise from simply using a HMM in our application. One is substantive: a key interest is to describe how career behavior varies across individuals, groups, and time; thus we naturally need to include covariates in the models. Another is more technical and concerns traces. A key component of the problem is that it is essential that the trace-emission process be well modeled: otherwise, errors here will be communicated to the rest of the procedure, which will result both in a poor reconstruction of X and in biased inference of the $(\beta^{x,x^{\prime}})$. We present for now the substantive reason for why it is essential to have covariates and lagged values of traces when modeling the trace-emission process. We defer examples based on synthetic data to section 3.2 as they require more notation.
Two problems make the inclusion of covariates and lagged values essential in modeling the trace-emission process: heterogeneity and self-excitation. In all the examples cited above, we expect heterogeneity between individuals emitting traces. Some journalists and academics write more articles than others, some civil servants make more important decisions and are promoted more often than their peers, and some athletes participate in more competitions than their competitors. If we do not account for this, our inference will confound the trace frequency variety and the actual behavior of careers. Furthermore, in most of the examples we cited, it is expected that some traces will be tied or will lead to other traces in quick succession, a fact that can either be described as self-excitation or time heterogeneity. Politicians writing political posts will lead to more posting, civil servants joining a committee will lead to more announcements when the committee arrives at a position, and journalists writing a piece on some hot-button issue can be expected to write more in the following weeks or months. Not accounting for that will lead to overestimating the probability of finding traces at any given time, and thus to overestimating the number of transitions across states of trajectories X. For this reason, using a simple HMM is not sufficient, and there is a need for both covariates and transformations of lagged values of traces.
On the other hand, it might be argued that a more complex model is called for. We resist the urge to go for more complex models, because Markov switching models exhibit properties that greatly facilitate the inference and enable us to efficiently target the posterior in a Bayesian framework. In particular, their factorization structure still enables the usage of the Baum-Welch algorithm when reconstructing sequences X from traces Y. One can take any presentation of HMM reconstruction (e.g. Marin and Robert, 2014), simply subscript the transition and emission rates, and perform the reconstruction unchanged. In fact, this would even work for rates that are obtained with more complex models than a GLM, such as neural networks (Tran et al., 2016), as long as the independence structure described in Figure 1 holds. We do not consider such extensions here, and remain in the GLM realm as our goal is to test substantive hypotheses.
This last point connects to one final discussion of the goals of the analysis. If the goal were only to reconstruct the trajectories X with the highest accuracy at the individual level, other solutions from (Bayesian) machine learning would probably be better suited, using recent developments in deep learning for sequences and inverse problems. However, our primary goal here is to provide a method that enables the testing of substantive research hypotheses, reconstruction being only secondary and used in an MCMC sampler. We still report quantities related to reconstruction accuracy with cross-validation experiments, but our main interest throughout is to enable the unbiased and precise inference of substantively motivated parameters. Yet, more complex developments could be motivated for the trace emission part (since we consider emission parameters $\eta^{x}$ to be nuisance), and for the career trajectory part (notably using varying-length Markov models, see e.g. Le Picard and Rossi, 2023).
2.2.4 MCMC inference sampler
We propose a tailored estimation scheme by MCMC to target the posterior, consisting of one Gibbs step for generating the missing values of X, and Metropolis-within-Gibbs steps for exploring the parameter spaces. This contrasts with a recent work from Michelot (2022) which proposes a faster (but non-exact) estimation using the Laplace approximation of the posterior. We give here a general presentation on our algorithm, but note that our implementation (available at https://github.com/theovoldoire/Hidden_Careers_Public) is more problem-specific: in our implementation, we only consider a state space of dimension 2, and perform several simplifications similar to those discussed in section 2.2.2 and introduced in section 3. The extension to the general case would be straightforward.
Gibbs step for unobserved values of $X$. The Gibbs step consists of simulating missing values of X given observed values of X, traces Y, and parameters $(\beta^{x,x^{\prime}})_{x,x^{\prime}\in\mathcal{X}}$, $(\eta^{x})_{x\in\mathcal{X}}$, and consists of two stages. In the first stage, we compute the transition parameters $\gamma_{i,t}^{x,x^{\prime}}$ for all individuals $i$, time $t$, and states $(x,x^{\prime})\in\mathcal{X}\times\mathcal{X}$; and the conditional probabilities $p(Y_{i,t}|X_{i,t}=x,\widetilde{W}_{i,t},\eta^{x})$ provided by the expected value $\lambda_{i,t}=\mathbb{E}[Y_{i,t}|X_{i,t}=x,\widetilde{W}_{i,t},\eta^{x}]=g(\widetilde{W}_{i,t}^{T}\eta^{x})$, for all $i,t$ and states $x$. In the second stage, we use the Baum-Welch algorithm. Our implementation is based on the R package HMM (Himmelmann, 2022), for which we simply added a time subscript on parameters; note that the Baum-Welch algorithm allows for transition and emission parameters which are not constant across time and units.
Metropolis-within-Gibbs step for parameters. We take inspiration from Marin and Robert (2014) and apply the same method successively to different parameters of interest. Iterating on $x\in\mathcal{X}$, we update separately each vector of parameters, first $(\beta^{x,x_{1}},...,\beta^{x,x_{K}})$ jointly, and then $\eta^{x}$. For each update of an arbitrary parameter vector $\mu$, we proceed as follows: compute the MLE $\hat{\mu}$ for this given parameter using either equation 3 or 4 and compute the associated covariance matrix $\hat{\Sigma}$. Second, run for $M_{\text{metropolis}}$ iterations a Metropolis-Hastings random walk with the proposition $\tilde{\mu}\sim\mathcal{N}(\mu,\tau^{2}\hat{\Sigma})$, with $\tau$ a hyperparameter that we set to 1 after experimentation. After $M_{\text{metropolis}}$ iterations, keep the final value of parameter $\mu$ as the new value. In our applications, visual checks indicated that it was sufficient to take $M_{\text{metropolis}}=30$.
3 Public-Private paths in French elite civil servants
In this second part of the article, we showcase the general methodology on real data taken from the social and political sciences: public-private paths of civil servants in France. While our focus is on the inferential process itself, we also discuss the applied literature and extend it in several manners.
3.1 Presentation of the studies
3.1.1 General framing
Paths of civil servants in the private sector in France have been studied by multiple authors (Bouzidi et al., 2010; Kolopp, 2021; Rouban, 2014) and have been a source of interest in the general public. We explore two aspects assessed as interesting in the literature but not yet described, or not described with such precision: how the structure of public-private paths has evolved since 1990 (Bouzidi et al., 2010); and how patterns of these paths vary across many organizations (Charle, 1987) or individual characteristics like gender (Kolopp, 2021). To this end, we will study two populations of elite civil servants: former students of the National School of Administration (École Nationale d’Administration, hereafter ENA) in Section 3.3, and members of powerful services (listed below) in Section 3.4. Previous studies have focused on detailed data sources only: well described career paths for a small number of people. As such, they have not explored thoroughly these questions, which require large amounts of data, and very diverse data.
We review the relevant applied literature in sections 3.3 and 3.4. We first introduce our data sources along some particularities of the analysis. Details for how we constructed and assembled the population can be found in Appendix A.
3.1.2 Detailed and trace data sources
Recall that we combine two data sources: a detailed data set for a smaller number of individuals, and a trace data set which covers the entire population but with less detail. Quantitative details on the size of each population and their coverage in the detailed source are provided in sections 3.3.1, 3.4.1 and in Appendix B.
Our detailed data come from the website LinkedIn. LinkedIn is a social media platform oriented towards professional networking, where users describe their past work experience on their profile. Many managers, both in the public and private sectors, have a profile, but not the entirety of the population. In addition, as these data are auto-biographical, individuals may choose to not include all relevant information: there are gaps in profiles, which means that there might be missing data even for individuals with a LinkedIn profile. This means that even when an individual is in this detailed data source, it can still be useful to complement the analysis with a trace data source.
Our trace data source is the Official Journal of the French Republic (hereafter JO or “Journal Officiel”), an administrative record with a significant legal value that includes new pieces of regulation, executive decisions and nominations. This source documents nominations for many top-level positions in public organizations. It also includes, in certain cases, information about promotions, other career changes (such as retirement) and decisions taken by authorities that might involve named civil servants. Importantly, certain key positions are systematically described, and this source can thus be used to define exhaustive and representative populations. However, not all public positions are described, and there is often no duration attached to these positions, which explains why it may only be treated as a traces dataset. The source was digitized starting in 1990, and this defines our temporal frame (1990-2022).
3.1.3 Specificities in the implementation
Indices and span of the analysis. We consider the trajectories of individuals through time $t=1,...,T$, where an increment by one unit corresponds to 6 months: $t=1$ corresponds to January-June 1990, and $T=68$ corresponds to July-December 2022. Each individual $i$ is modeled on an interval $[T_{min}^{i},T_{\max}^{i}]$. We let $T_{\min}^{i}$ be their first membership-defining trace (see Appendix A). If the individual has a trace that states they retired, we use that date for $T_{\max}^{i}$; else we let $T_{\max}^{i}=T$, where $T$ corresponds to our last observation time (July-December 2022). Individuals are thus modeled on different spans, a fact that is possible in the general framework but makes notation more cumbersome.
State spaces and parameters. As our aim is to describe public-private paths, we study the state space $\mathcal{X}=\{0,1\}$, with 00 corresponding to the public sector and $1$ to the private sector. It is possible to describe heterogeneity of the public sector by including information in the covariates W and without increasing the size of the state space. We thus have two parameters of interest indexed by individuals and time: $\gamma_{i,t}^{0}:=\gamma_{i,t}^{0\rightarrow 1}$ the probability to exit the public sector at time $t$, and $\gamma_{i,t}^{1}:=\gamma_{i,t}^{1\rightarrow 0}$ the probability for an individual who had previously left to the private sector to return to the public sector. Finally, we describe traces using simply a dichotomous variable as this led to better results: we take $\mathcal{Y}=\{0,1\}$, with $Y_{i,t}=1$ if we observe at least one trace in JO in the given six-months interval and $Y_{i,t}=0$ otherwise. This is modeled as $Y_{i,t}\sim\text{Bernoulli}(\lambda_{i,t})$.
Configuration for autoregressive traces. We code three variables to describe the autoregressive aspect of traces. Two are given by $A_{i,t}(\varphi)=({\sum_{0<s\leq t}\varphi^{s}Y_{t-s}^{i}})/({\sum_{0<s\leq t}\varphi^{s}})$, where we set $\varphi=1$ to prioritize long-term information and $\varphi=0.8$ for medium-term information. $A_{i,t}(1)$ and $A_{i,t}(0.8)$ take values between 00 and $1$. For short-term information, we create a transformation of the time lapsed since the last trace
$L_{i,t}=\sqrt{\max\{p\ |\ \forall s<p,Y_{t-s}=0\ \text{and}\ Y_{t-p}=1\}-1}.$ (5)
These three variables provide a low-dimensional and efficient encoding of past traces and tackle the issue of the varying size of past values, but many other encodings could be motivated since these are not our parameters of interest. We used the cross-validation reconstruction loss criterion introduced in section 3.3.3 to evaluate different possible values for $\varphi$ (we tried $\varphi=0.6,0.7,0.8,0.9$) and found that $\varphi=0.8$ minimized the loss.
Handling of boundary effects. An issue is that when $t$ is low, there are not many past trace variables. This becomes problematic because a specificity of our setting is that we use traces both to define the population (as the starting positions we are interested in are systematically described) and as a complementary source (for all traces). This means that by construction, $Y_{i,T^{i}_{\min}}=1$, which needs to be taken into account. As a solution, we ignore this first trace (setting the value to zero), and include specific covariates to handle points before any traces were found: $\textbf{1}_{\{A_{t,i}(1)=0\}}$, and $L_{t,i}\times\textbf{1}_{\{A_{t,i}(1)=0\}}$. A more complex handling is required when including time-sensitive coefficients, and we include instead an interaction between B-splines for age and the indicator $\textbf{1}_{\{A_{t,i}(1)=0\}}$ (see below, models 3, 4, 5 of Study 1).
Prior distributions and convergence. Results presented throughout are with non-informative, flat (improper) prior distributions. Other tests were performed with low-informative priors, e.g. $\mathcal{N}(0,\sigma^{2})$ with $\sigma=5$ for model 3 in study 1, with unchanged results. Contrary to fully unsupervised methods where the dependency on the prior choice could be higher, our setting only exhibits a small dependency, as points taken from the detailed data source are not simulated an thus reduce the relative importance played by the prior in the posterior distribution. We run our MCMC sampler on 5000 iterations and discard the first 200 as a burn-in period. Visual checks show that this is sufficient to guarantee convergence and the MCMC sampler mixes well. Throughout studies 1 and 2, the ESS of all parameters is above 200, and the $\hat{R}$ below $1.006$, indicating that the Markov chain has converged and has appropriately explored the posterior distribution.
3.2 Empirical demonstrations on synthetic data
We present empirical demonstrations of multiple claims made in Section 2, taking advantage of the notation and intuitions that arise from this specific case. We aim to demonstrate two claims: in a reasonable case with a well-specified model, researchers can expect to gain substantial statistical power by adopting our procedure; and when using a standard HMM despite having heterogeneity in the trace-emitting process, errors there communicate and lead to erroneous conclusions for interest parameters, further motivating our approach.
Gains of statistical power. For the first claim, we simulate trajectories using the following form for parameters, with $g^{-1}$ the inverse logit transform:
$\forall i,t,\ \gamma^{0\rightarrow 1}_{i,t}=g^{-1}(-4+C_{i}),\ \gamma^{1\rightarrow 0}_{i,t}=g^{-1}(-4),\ \lambda_{i,t}=g^{-1}(-1.5),$ (6)
with $C_{i}\sim\mathcal{U}([0,1])$ an individual observed variable, that could correspond in our example to the cohort of individual $i$. Our main parameter of interest is the impact of the cohort on $\gamma_{0}$. These coefficients are similar to estimates provided for the simplest models in Section 3.3, so this simulation experiment can be deemed realistic. Each trajectory is initiated at state 00, and is of length 60. We simulated three datasets, with $N=200,500,$ and $1000$ individuals, where everyone is observed in the first two cases (and so no data augmentation is performed), but only 200 individuals in the third case.
Table 1: Reporting on the first synthetic data example
Case 1 Case 2 Case 3 True value
n 200 500 1000
$n_{\text{obs}}$ 200 500 200
$\gamma^{0}$, Intercept [-4.23,-3.52] [-4.25,-3.81] [-4.2,-3.81] -4
$\gamma^{0}$, Cohort [-1.05,1.25] [0.16,1.61] [0.74,1.98] 1
$\gamma^{1}$ [-4.92,-4.05] [-4.1,-3.69] [-4.34,-3.84] -4
$\lambda$ [-1.54,-1.41] [-1.52,-1.44] [-1.5,-1.44] -1.5
We report on the concentration of the posterior of coefficients, by reporting on 90% credibility intervals in Table LABEL:synth:ex1, from which different conclusions can be drawn. First, case 3 performs better than case 1, especially for the cohort coefficient for $\gamma^{0}$, where the posterior is much more concentrated on the true value. Second, cases 2 and 3 can be said to have relatively similar performances, with comparable posteriors and uncertainty. However, the gain here is that instead of observing 500 individuals, only 200 individuals had to be observed in the detailed data source, and imprecise knowledge using traces compensated precise knowledge on a population more than twice larger. The exact population size numbers will vary from application to application, the main idea being that the more “informative” the traces are, the more useful the data augmentation procedure will turn out to be.
Issues of standard HMM misspecification. For the second claim, we simulate trajectories using the following form for parameters
$\forall i,t,\ \gamma^{0\rightarrow 1}_{i,t}=g^{-1}\left(-4+B_{i,t}\right),\ \gamma^{1\rightarrow 0}_{i,t}=g^{-1}(-4),\ \lambda_{i,t}=g^{-1}(-2+2.5A_{i,t}(0.8)),$ (7)
where $B_{i,t}=\frac{t}{60}$ is a time-varying covariate, which here could correspond to age or period, and $A_{i,t}(0.8)$ the autocorrelation covariate as defined earlier. We simulate 2000 sequences of length 60 and observe 30% of them. We implement two models: one well-specified (ours), and one misspecified (standard HMM), where covariate $A_{i,t}$ is not included.
Table 2: Reporting on the second synthetic data example
Well-specified Misspecified True value
$\gamma^{0}$, Intercept [-4.28,-3.98] [-3.74,-3.49] -4
$\gamma^{0}$, Age [0.86,2.02] [-0.3,0.7] 1
$\gamma^{1}$, Intercept [-4.35,-3.95] [-4.1,-3.79] -4
$\lambda$, Intercept [-2.03,-1.98] [-1.49,-1.45] -2
$\lambda$, $A(0.8)$ [2.4,2.6] 2.5
Coefficients are reported in Table LABEL:synth:ex2. In the well-specified case, the output credibility intervals correctly contain the true parameter values. On the other hand, in the misspecified HMM case, the coefficient for the age effect $B_{i,t}$ is wrong in sign and magnitude. Such problems will arise for time-varying covariates when an autocorrelation behavior is not taken into account in the modeling phase. The same point is also shown by the cross-validation reconstruction losses obtained in Study 1 below.
3.3 Study 1. ENA graduates and temporary exit
3.3.1 Motivation and population definition
The National School of Administration ENA was founded in 1945 to select and train high civil servants for the French state. Graduates of this school (around 80 to 100 a year) play a central role in the political and administrative life of the country, and they constitute a natural population when studying elite civil servants (Bouzidi et al., 2010; Rouban, 2014). We study students admitted between 1990 and 2019, before the school was replaced by the National Institute for Public Service (INSP) in 2021111Data taken from Journal Officiel were first assembled and cleaned up by Nathann Cohen and put up on an API on the website steinertriples. Throughout the article, we rely on the amazing meta-data provided by this API, in addition to our own data cleaning procedures.. We only keep track of individuals who graduated from ENA and whose first position was working for the state, which we operationalize by only keeping individuals for which at least one trace can be found in the three years following their graduation. The number of students identified using this method by year of admission can be found in Appendix B and forms an exhaustive total of $N=2715$.
ENA graduates constitute a natural population when studying public-private career paths, as they hold a central position in the organization of the French state. Rouban (2014) conducted a survey over 6 student year-groups ($N=620$) who graduated in 1969-1970, 1989-90, and 1999-2000. He reports that, 10 years after graduating, 8.3% of students from the first two cohorts were working in a private company, compared to 14.5% for the two next cohorts, and 9.5% for the last two, which indicates year-to-year variation but not a clear trend towards increasing exit rate to the private sector.
The results of Rouban (2014) seem to show that the average time taken for an ENA graduate before leaving the public sector has drastically decreased over time, which would imply a stark increase in the exit rate from the state. We must however note that their methodology does not allow for comparing different cohorts. Indeed, many statistics considered in this study, like average time to leave to the private sector, are subject to right-censoring; this was not accounted for or acknowledged. This is problematic as the observation frames are significantly different across cohorts: it was, at the time of publication, 44 years for the 1970 cohort, but only 14 years for the 2000 cohort. We assume the difference with our results is due to our methodology properly taking into account the different lengths of observation frames. We conclude that the variation reported by Rouban (2014) is artefactual. The other quantitative conclusions of Rouban (2014) are not subject to this bias, and are coherent with our results.
Bouzidi et al. (2010) focus on a subgroup of ENA graduates, of the “civil administrator” status working in the Finance Ministry, between 1960 and 2002 ($N=438$), using hand-collected data. They show that, on their sample, about 40% of individuals went at least once into the private sector 20 years after graduating. The advantage of their approach is that they used a formal duration analysis model and are not susceptible to right-censoring biases. They however do not attempt to estimate whether there was an acceleration or not on their period of study, and stress that doing such inference is difficult for later cohorts because of right-censoring. They also estimate time spent in the private sector using simulations using two Weilbull distributions, one for the public sector and one for the private sector. Our work can be seen as a discretization of such a procedure that also enables missing data imputation and inclusion of covariates.
We build our first study from these surveys and provide two improvements. First, thanks to our methodology, we can broaden the scope and analyze all students who graduated between 1991 and 2021, including the 55% do not have a LinkedIn profile, as well as the individuals who have a partially incomplete LinkedIn profile. In addition, since our modeling accounts allows individuals to leave for the private sector and then return to the public sector, we can measure more precisely the public exit rate, and distinguish between individuals the public sector permanently or temporarily.
3.3.2 Models
We implement four initial configurations which are described here from simplest to most complex. Models 1 and 2 are introduced for pedagogical purposes. Model 1 is a standard hidden Markov model: the covariate vector $W_{i,t}$ only includes the intercept. Model 2 is a Markov switching model for which only the autocorrelation covariates described in section 3.1.3 are included for the estimation of $\lambda$.
Models 3, 4 (and 5) include time-specific interest covariates and notably age. Age is encoded as time elapsed since the individual was admitted to ENA, on a range from 0 to 1 (with 1 corresponding to the entire time frame, 32 years). Age is modeled in a semi-parametric fashion, using B-splines of degree 2, with 5 degrees of freedom including an intercept, and boundary knots in 0 and 1, leading to $g=1,...,5$ covariates $S_{i,t}^{g}$ in the linear model (with the R package splines). We do this in order to saturate the model on the age dimension, so we can safely identify an effect related to cohort or period effects. Using splines works better than multiple powers of age: under this parameterization, parameters are less correlated, so the sampler mixes faster. As the spline basis already includes an intercept, no general intercept is included. Recall from our discussion in Section 3.1.3 that we have to use a more complex scheme to handle boundary effects for $\lambda$, and interact these five covariates with $\mathbb{I}_{A(1)=0}$ for all three models. More specifically, Model 3 adds a time-varying component depending on the age of individuals for $\gamma_{0}$ and $\gamma_{1}$. Model 4 adds a cohort effect for the estimation of $\gamma^{0}$, $\gamma^{1}$, and $\lambda$. Cohort is encoded as the time at which the individual was admitted at ENA, ranging between 0 and 1.
Finally, we showcase that it is possible to answer questions that are less descriptive and more centered on substantive social mechanisms by studying the dependence between public-private paths and electoral cycles. We implement an additional Model 5, which is similar to Model 4, except for the fact that instead of a cohort effect, we estimate a period effect using a dummy variable $P_{i,t}$ with value $1$ if $t$ or $t+1$ is an election time point, and 00 else. Election time points correspond to Jan-Jun 1995, 2002, 2007, 2012, 2017, 2022 (French presidential elections). This dummy is shifted by one unit (so $t+1$ or $t+2$) for $\lambda$.
3.3.3 Analysis of Model Accuracy
We first analyse the output for models 1, 2, and 3. We study both the posterior of parameters for simpler models and the quality of data augmentation by performing analyses where detailed data source information is left out for some observed individuals.
Parameters posterior. The simplicity of models 1 and 2 enables full reporting on the posterior of their coefficients, which we do in Table 3 with their 0.05, 0.50 and 0.95 quantiles. Full information about model coefficients can be found in Appendix D.
model $M_{1}$ $M_{2}$
quantile q05 q50 q95 q05 q50 q95
$\gamma_{0}$, (Intercept) -4.233 -4.171 -4.111 -4.621 -4.542 -4.465
$\gamma_{1}$, (Intercept) -4.054 -3.938 -3.83 -3.667 -3.525 -3.391
$\lambda$, (Intercept) -1.51 -1.492 -1.474 -1.853 -1.786 -1.718
$\lambda$, $L$0.16 -0.134 -0.108
$\lambda$, $A(1)$0.133 0.088 0.315
$\lambda$, $A(0.8)$ 2.027 2.249 2.467
$\lambda$, $I(A(1)=0)$ 0.102 0.191 0.279
$\lambda$, $I(A(1)=0):L$0.173 -0.137 -0.101
Table 3: Coefficients for Models 1 and 2.
Comparison between models 1 and 2 shows why there is a need for autoregressive coefficients and to handle boundary effects, as discussed in section 3.1.3. The main effect of adding autocorrelation coefficients from Model 1 to Model 2 can be observed in the difference between their median intercept for $\lambda$. For Model 1, the median intercept is at $-1.492$ ($I_{90}=[-1.51,-1.474]$), which essentially corresponds to an average probability of 0.22 for an individual to generate a trace at any point in time, which is higher than the one obtained for Model 2, at $-1.786$ ($I_{90}=[-1.853,-1.718]$), and which is more in line with what could be expected. All credible regions for autocorrelation coefficients except $A(1)$ do not contain zero despite that they are quite correlated by construction. Finally, the coefficient responsible for handling the boundary effect $\mathbb{I}_{A(1)=0}$ is very different from zero and negative, showing indeed that it is needed. Configurations without it also have very different values for all other coefficients.
Evaluation by reconstruction loss. We repeat the inference under models 1, 2, 3 but with some data withheld. In this experiment, we exclude from the data 20% of individuals in the detailed data source. We can thus assess the data augmentation quality could on points $(i,t)\in S_{\text{obs.,exc.}}$ that are observed but excluded from the analysis. We adopt a simple cross-entropy loss on the trajectory for such points $(i,t)\in S_{\text{obs.,exc.}}$, $l_{i,t}=-(X_{i,t}^{\text{obs}}\log(X_{i,t}^{\text{pred}})+(1-X_{i,t}^{\text{obs}})\log(1-X_{i,t}^{\text{pred}})$), where predictions are computed as the average of observations generated in the Gibbs step over the last $300$ iterations of the chain, with probability clipping at 0.01 and 0.99, and observations are taken from the detailed data source. For Model 1, we also implement a version where no information about the traces was used at all, which we denote Model 1b. Full reporting can be found in Appendix D.
The average cross-entropy loss obtained for the different models is 0.39 for Model 1, 0.29 for Model 1b, 0.22 for Model 2, and 0.21 for Model 3. A first very clear conclusion is that HMMs without autoregressive components are not suited for the task, as Model 1b – which did not even have access to traces! – performed better data reconstruction than Model 1. We observe however that better reconstruction is obtained once including autoregressive components (in models 2 and 3). The change in accuracy from Model 2 to Model 3 is small in size, but our interest is rather to perform hypothesis testing when including the additional age coefficients. Overall, this indicates that one could obtain better reconstruction by adopting more complex models, but this would be at the cost of model interpretability and hypothesis testing, which is our main focus.
3.3.4 Substantive Results
Figure 2: Median marginal probabilities across time given in Model 3. $\gamma^{0}$ refers to the probability of leaving the public sector, and $\gamma^{1}$ the probability of coming back.
Dependency with age and cohort. Figure 2 summarises information related to the probability of transitioning between states depending on age for model 3. It was computed by taking the inverse logistic transform of the sum of the intercept with the age spline coefficients. We first observe, that, around 0 years after admission, the probability of leaving for the private sector $\gamma^{0}$ is very low; this was to be expected, since individuals who never leave any trace (e.g. because they immediately resigned from ENA) are not included in the dataset. The probability $\gamma^{1}$ that an individual moves back from the private to the public sector is very high, but note that this concerns a very slim proportion of the population (less than .$7\%.$ up to three years). The median probability $\gamma^{0}$ then increases to $0.013$ where it remains more or less constant up to 25 years after admission. A stabilization occurs for $\gamma^{1}$ at around double the rate, at $0.025$. This is interesting as we observe there isn’t an important variation in the probability of leaving across the career, and the hypothesis of an overall geometric rate for leaving and coming back is quite realistic. We interpret the change of behavior 30 years after admission as artefactual, as almost no one in the population has had the time to arrive at this stage, which can also be seen with the widening of uncertainty.
Figure 3: Posterior distribution of selected parameters for models 4 and 5. Left: effect of the cohort $\beta_{\text{cohort}}^{i}$. Right: effect of the indicator that it is an election year $\beta_{\text{year}}^{i}$. Top: Posterior of the parameter $\beta^{0}_{\text{coeff}}$ measuring the impact on the probability of leaving the public sector $\gamma^{0}$. Bottom: Posterior of the parameter $\beta^{1}_{\text{coeff}}$ measuring the impact on the probability of returning $\gamma^{1}$.
We now turn to our main quantities of interest, cohort-related coefficients, which are presented in the left pane of Figure 3. The view is quite contrasted between $\gamma^{0}$ and $\gamma^{1}$. On the one hand, it isn’t possible to conclude that the chance of leaving for the private sector $\gamma^{0}$ has significantly changed across the period, as the credible interval well includes zero ($\text{med}=-0.226,I_{90}=[-0.625,0.176]$). However, it is at least possible to state strongly that this coefficient has not increased across time, and has quite possibly decreased. On the other hand, we can state with fair confidence that $\gamma^{1}$ has increased over time, with $\text{med}=1.16$, ($I_{90}=[0.301,1.98]$). These results are robust to misspecification, as shown by checks with modular inference described in Appendix C. We thus conclude that on the one hand, the probability of leaving the public sector has not increased in the period, and may even have decreased; and that the probability of coming back has increased in the period, contrary to what we expected before conducting this study.
Dependency with the electoral process. The posterior for the election time effects are reported in the right pane of Figure 3. We observe that the overall size of potential effects is far smaller than for period effects, and so should be interpreted cautiously. On the one hand, we conclude that elections have no impact on $\gamma^{0}$, as the credible interval is very well centered on zero ($\text{med}=-0.0212,I_{90}=[-0.249,0.201]$). On the other hand, however, despite important uncertainty, there seems to be a small positive effect of election time on $\gamma^{1}$ ($\text{med}=0.302,I_{90}=[0.000,0.302]$). In summary, electoral shifts probably do not influence the chance of high civil servants to leave for the private sector, but it might lead to higher chances of coming back for those who left.
3.3.5 Discussion of the substantive results
We conclude that exit rates from the state have not increased on the period for former ENA students, contrary to what we expected before conducting this study. If any change was to be inferred from data, it would rather be that elite civil servants have now a lower probability of leaving for the private sector than before, but it is possible to remain undecisive. Furthermore, the rate of former civil servants coming back to work in the public sector after leaving has significantly increased over the period. In other words, when civil servants leave for the private sector, they now have a higher chance of coming back. This is notably a contribution with respect to (Bouzidi et al., 2010), who stressed that they could not study such evolutions given the constraints of their data despite the demands of their journal referees, and showcases the advantages of studying larger groups via data augmentation.
In addition, if there seems to be an interaction between political majority changes with election years and revolving doors for this population, it seems to affect the chance of coming back from the private sector than the probability of leaving.
3.4 Study 2. Public-private paths across the state
Many social scientists have early insisted on the usefulness of adopting a comparative point of view when describing civil servants populations and public-private paths (Charle, 1987). We are able with our method to describe the heterogeneity of public-private paths across the State and how they interact with gender.
3.4.1 Population definition
Top administrators are divided between groups and services that are sometimes tied to certain special statutes. While recent regulatory changes have attempted to increase mobility between them and foster a common organizational culture, they remain useful categories to base our approach upon.
Our study focuses on the following groups: three inspection groups whose mission is to audit public organizations organized thematically around financial affairs (igf, Inspection Générale des Finances), social affairs (igas, Inspection Générale des Affaires Sociales), and administration (iga, Inspection Générale de l’Administration); the Cour des Comptes (ccomptes), the supreme financial jurisdiction with a similar financial auditing function; the Conseil d’État (ce), whose members serve both as judges in the Supreme Administrative Court and as legal advisors to the government; the préfets (cprefet), state administration representatives in regions and departments; ambassadors (cdiplo, Corps Diplomatique); administrators of the national statistical and economics office (insee) who can serve key roles in the definition of economic policies; and finally, directors and assistant directors of two powerful services of the Ministry of the Economy, treasury (dgtresor), and public finances (dgfip).
This selection is voluntarily diverse and encompassing, as one modeling goal is to describe how public-private mobility is a heterogeneous phenomenon. We describe in Appendix A the method we use to link individuals to each group. The total number of individuals in each group alongside the proportion of them matched to a digital profile is given in Appendix B. Note that some individuals are affiliated with multiple services because their affiliation may change over time, a technicality we discuss below. The number of missing values is greater than in study 1, reaching 75%, which makes data augmentation even more valuable.
3.4.2 Relevant applied literature and hypotheses
Different researchers have studied the trajectories of high civil servants in France by defining populations based on organizational belonging at a certain time point, especially by focusing on their past trajectory.
Rouban (2010) studied the finance general inspection (igf) on the 1958-2008 period. igf appears to have an important place in the revolving doors phenomenon, as they identified that over the 578 inspectors present in the organization between 1958-2009, a majority had wandered at least once in the private sector (62%). They identify a discontinuity around 1990, with a massification of individuals having done a business school before attending ENA and then igf. They also observed after 2000 a growing proportion of igf members who have had an experience in the private sector before coming back, which seems quite aligned with results obtained in our first study with a positive cohort effect for $\gamma^{1}$. However, we underlie that this study too is subject to bias because of the lack of accounting of right-censoring.
Kolopp (2021) conducted a study on the general direction of the treasury (dgtresor) in the 1965-2010 period. They exhibited a gendered behavior: men are more likely than women to leave for the most prestigious private industries (especially banking); on the other hand, they do not seem to show gender differences in the overall probability of leaving the public sector. They describe that differentiated practices generate unequal outcomes at different points in careers, for example with the stress on international mobility for career advancement and differences in professional networks. If we have reasons to assume these factors also play a significant role for other organizations, their strength may vary, and other factors might come into play. Even if we don’t differentiate within private sector activities, we will compare behavior between men and women across the different organizations of our sample.
We attempt to build upon their results and propose a complete and systematic comparison across the different organizations we mentioned. This has value in the sense that it can qualify statements regarding the most prestigious organizations that are represented in these studies: are they really that special compared to other powerful services? For example, are the Finance Inspection and Treasury that much more affected by revolving doors compared to other services in the economy ministry or other inspections? We can also test statements linked to individual properties: do men have higher chances to leave for the private sector, and then come back?
3.4.3 Models
We encode organizational membership using a set of $D$ dichotomous variables $M_{t,1}^{i},...,M_{t,D}^{i}$ corresponding to our $D=11$ selected organizations, that we will include in the covariate matrices $W_{t}^{i}$. Contrary to the first study, we need to account for the fact that individuals can be affiliated to multiple organizations during their professional career. To remain consistent with our choice to focus on the main organizational identity of individuals, we impose that someone, at any given time, may only be linked to one organization in our set of $D$ organizations. This means we impose $\sum_{d=1}^{D}M_{t,d}^{i}=1$. If an individual is observed to become part of organization $d_{1}$ at time $t_{1}$ and then of organization $d_{2}$ at time $t_{2}$, we will have $M_{t,{d_{1}}}=1$ for $t_{1}\leq t<t_{2}$, and $M_{t,d_{2}}=1$ for $t\geq t_{2}$. This is a simplification, which has the advantage of enabling an analysis that is closest to what we could obtain by performing completely separate analyses for the different public organizations, even if it could be altered to study how having been the member of a given organization can affect behavior once being a member of another one included in the study.
We implement two configurations. Model 1 is a Markov switching model for which we only include an intercept for each group (no general intercept, and no reference organization), in addition to the autoregressive covariates for traces. The autoregressive covariates are interacted with the dummy of every organization. This is our reference model. Model 2 adds to the reference model a dichotomous variable encoding gender ($G^{i}=1$ if $i$ is a woman), for both $\gamma^{0},\gamma^{1}$ and $\lambda$.
3.4.4 Results
Reconstruction loss. The reconstruction loss obtained by the two models in this study is of $0.279$ and $0.265$, which is higher than in study 1. This can be explained because the inference task is more complex given the higher heterogeneity of the population, and is in line with what could be expected given the nature of the application.
Comparison across organizations. We report the $5,50,$ and $95$ quantiles of the posterior for each of these models. Median estimates of coefficients for Model 1 are reported in Figure 4, alongside their 90% credible interval, expressed in raw probability. Bear in mind that these are average coefficients that do not correspond to local probabilities individuals may have to leave at any given point in time, which are modulated by the observed traces and the local $\lambda$’s. As coefficients for dgfip stand out from the rest in both estimate and uncertainty, we do not report them in the Figure but do so in the text below.
Figure 4: Median and 90% credible interval of coefficients $\gamma^{0}_{\text{group}}$, $\gamma^{1}_{\text{group}}$ for individuals belonging to groups defined in subsection 3.4.1. Coefficients may be interpreted as average exit ($\gamma^{0}$) or come-back ($\gamma^{1}$) rates, yet it does not mean the model predicts transitions to happen at a constant rate.
We can firstly observe and rank organizations depending on the chance individuals have to leave for the private sector when they are part of them. On the side of organizations with the highest departure rate, we find the General Inspection of Finances ($\gamma^{0}_{\text{igf}}=0.031$) and Prefects group ($\gamma^{0}_{\text{cprefet}}=0.029$). In an intermediary situation, we find the Diplomats group ($\gamma^{0}_{\text{cdiplo}}=0.022$), and the General Inspection of Administration ($\gamma^{0}_{\text{iga}}=0.020$) At the lower hand of the spectrum, we find the statistical office ($\gamma^{0}_{\text{insee}}=0.016$), the Cour des Comptes ($\gamma^{0}_{\text{ccomptes}}=0.015$), Treasury ($\gamma^{0}_{\text{dgtresor}}=0.014$), the General Inspection for Social Affairs ($\gamma^{0}_{\text{igas}}=0.013$), and the Conseil d’État ($\gamma^{0}_{\text{ce}}=0.011$). Not shown in the table, the General Direction for Public Finances has the lowest probability ($\gamma^{0}_{\text{dgfip}}=0.006,I_{90}=[0.004,0.009]$). This last very low coefficient means that there are almost no departures from this direction, and when that is the case, civil servants move first to another service included in this study. The overall ranking should be intuitive to the reader accustomed to the question – the model rightly identifies that individuals in the Treasury have a higher chance of leaving for the private sector than individuals in the Inspection for Social Affairs. A seemingly counterintuitive result is the position of igf and ccomptes: these two groups have a similar function (financial auditing of the state), but have very different positions in Figure 4; this observation is in fact aligned with an effect identified in the survey from Charle (1987).
We observe interesting contrasts in the chance of individuals to come back working in public administration after being in the private sector. Inspection of Finances again clearly stand out, with ($\gamma^{1}_{\text{igf}}=0.024$). With varying levels of uncertainty, Conseil d’État ($\gamma^{1}_{\text{ce}}=0.020$), Inspection of social affairs ($\gamma^{1}_{\text{igas}}=0.019$), Treasury ($\gamma^{1}_{\text{dgtresor}}=0.018$), and Cour des comptes ($\gamma^{1}_{\text{dgtresor}}=0.016$), they appear in an intermediate situation. At the lower end of the spectrum, we observe the statistical office ($\gamma^{1}_{\text{insee}}=0.006$), inspection of administration ($\gamma^{1}_{\text{iga}}=0.004$), prefects ($\gamma^{1}_{\text{cprefets}}=0.003$), and diplomats ($\gamma^{1}_{\text{insee}}=0.002$).
We thus conclude that the two rankings are not convergent and would even be negatively correlated if we excluded igf. However, we find it more accurate to describe this as forming three groups. First, igf constitutes a group on its own, standing out from the rest as having both high probabilities for leaving and coming back. Classical results claiming that moving into the private sector is part of the career advancement process hold well, and we further prove that this is quite distinctive to this organization. We identify a second group of organizations where the chance of leaving is relatively high, but the chance of coming back is rather low: this concerns the statistical office, the inspection of administration, diplomats and préfets. The situation is unsurprising for the first two, which are known to be closer to the private sector than other organizations. For préfets and diplomats, this might be because we study individuals who are simultaneously at the top of certain services but without as high position security as in other organizations (such as Conseil d’État, where people can keep their status): people (involuntarily) leaving could find themselves unable to come back. Finally, we find a group of organizations where the chance to leave is comparatively low and the chance to come back high: Conseil d’État, Inspection of Social Affairs, Cour des Comptes. One could also place Treasury in this group, but parameter uncertainty (especially on $\gamma^{1}$) requires caution. This concerns organizations for which members acquire life-long status that can help them come back after leaving; and also for which clear career advancement schemes exist in the public sector.
Interaction with gender
The posterior of the odds ratio for women versus men can be found in Figure 6 as a way to summarise interest parameters for model 2. A distribution on the right means we have grounds to think women have a higher chance of leaving for the private sector than men, and vice-versa. A distribution centered on 1 indicates we should remain indecisive on the matter.
Figure 5: Posterior distribution of $\frac{\gamma^{0}_{\text{group},W}/(1-\gamma^{0}_{\text{group},W})}{\gamma^{0}_{\text{group},M}/(1-\gamma^{0}_{\text{group},M})}$, the odds-ratio of the differential of exit rate between women and men expressed, for each of the nine groups in Study 2; the acronyms of the groups are defined in subsection 3.4.1. When the posterior is concentrated on the left-hand side, we can conclude that women in that group have a lower probability than men of exiting the public sector.
Not all organizations have the same profile of differentiated practices between men and women, but a broad conclusion can nonetheless be inferred: women have less chance than men to leave for the private sector. For certain organizations, the difference is strong and identifiable: Cour des Comptes, Inspection of Administration, Inspection of Social Affairs (odds ratios below 0.7). The difference is smaller but still observable for Inspection of Finances, Préfets, and Diplomats. The situation is undecided for Conseil État and the Statistical Office, the latter suggesting a small positive effect on the side of women. Treasury here stands out in terms of uncertainty (as there are less than 200 individuals to work on and a small proportion of women), but the side of the effect would be rather positive for women.
3.5 Discussion on the substantive results
We find three profiles of organizations depending on the chance of leaving and the chance of coming back: one with high chances of leaving but low chances of coming back; one with low chances of leaving but high chances of coming back; and Inspection of Finances standing out from the rest with high chances of both to leave and to come back. Women have overall less chance to leave for the private sector than men, but the situation is not constant across public organizations.
4 General discussion
We have proposed a general methodology for studying career trajectories on large populations, relevant for many problems in the social sciences. This method relies on two data sources, one with high precision but low coverage, and another, consisting of imperfect traces, but with high coverage. Markov switching models with autoregressive components constitute a powerful and flexible tool that can be fruitfully used to analyze this kind of data. They can be well estimated using a custom MCMC sampler combining a Gibbs step and a Metropolis-within-Gibbs step, but remain flexible enough to test substantive research hypotheses. In addition, we showcased the issues of using misspecified HMMs using synthetic data examples and showed one could gain statistical power using our method.
We have showcased the usefulness of our approach in two applied studies on French civil servants and their trajectory into the private sector, extending the social scientific literature along the way. An innovation in our applied studies is found in the way data were collected and cleaned. Our philosophy was to automate as much work as possible, for example in the name-matching phase, which enabled us to study a population an order of magnitude larger than what is usually done in traditional elite research (prosopography studies rarely exceed a few hundred individuals). However, we were not able to automate everything away, like public-private organization name classification, and the main limitation would again become human time if a researcher attempted to scale up this work by another order of magnitude. Our MCMC sampling procedure is, compared to approximate calculations, more costly, which could also be a problem when trying to scale up this approach. We found that the main cost of the procedure was in the computation of the MLE in the initialization of each Metropolis step, in addition to the random walk which required multiple computations of the likelihood. The cost of the Gibbs step was less important as it was very easy to parallelize it across individuals. Enabling a better scaling of the procedure could enable to study of a larger population, at a finer scale than based on six-month increments.
Acknowledgments
The authors are grateful to four anonymous reviewers, an associate editor, and an editor for their very helpful comments on a previous version of this manuscript. RJR received funding from the European Union under the GA 101071601, through the 2023-2029 ERC Synergy grant OCEAN.
Appendix A Data assemblage
This appendix documents the various procedures we used to assemble the dataset. To assemble data, we queried all traces of the JO linked to the corresponding organizations and manually crafted a rule-based approach to only keep those that define membership. All traces and LinkedIn profiles are attached to a full name. There is therefore a high risk of data being tarnished by homonyms. In this section, we describe the careful procedure used to eliminate this risk as much as possible, while preserving information.
A.1 Selection of administrative traces
For every person in our population, we query the Journal Officiel to obtain every administrative order carrying the same name, which yields a first set of traces we name $S_{1}$. As many people have perfect homonyms, we have to filter this set. Most traces are tagged with one or multiple organizations, which range from small bureaus to the entire police force. To reduce this risk of perfect homonyms, we only keep traces linked to organizations significantly over-represented in $S_{1}$ compared to the entire Journal Officiel. For each organization with the set $O$ of traces linked to it, we compute:
$\frac{\mathbb{P}[O|S_{1}]}{\mathbb{P}[O]}=\frac{\#\{\text{traces linked to org}\ O\}\cap S_{1}}{\sum_{x}\#\{\text{traces linked to org}\ x\}\cap S_{1}}/\frac{\#\{\text{traces linked to org}\ O\}}{\sum_{x}\#\{\text{traces linked toorg}\ x\}}$
which can be interpreted as a relative risk score. We only keep traces linked to organizations with a score above 2, or above 1 if they have fewer than 1000 traces, or above 0.5 if they have fewer than 100 traces. This effectively removes organizations such as the army or the police where many promotions are described for the entire force, and constitute the main sources of perfect homonyms in $S_{1}$. For traces that are not linked to any organization, we also include them if they are a “délégation de signature” (meaning an authority delegates administrative power to the civil servant, a practice that remains limited in scope) or if it states an admission into a “grand corps”. This enables a good tradeoff between minimizing the homonymy risk and keeping as much data as possible, and yields a second set of traces $S_{2}$. For the first population, we then manually search for every individual for which we cannot find any trace upon graduation, and document the cases of name-changing for individuals who got married during their time as students (around 200 individuals) and add the relevant traces to the total after filtering them using the same procedure. For the second population, if a trace in $S_{2}$ documents the person has had another name, we also add the relevant traces after passing them through the same filters as before (this also represented around 200 individuals). This yields our final set of traces $S_{3}$.
A.2 Linkage with digital profiles
For each person in our population, we search on LinkedIn for up to three profiles that share the same name, download them, and then only keep the ones that list an experience in our starting set of organizations. For example, if an individual is included because they are part of the Statistical office, we only keep profiles with such an affiliation, using a rule-based approach. When more than one profile qualifies, we score the candidate profiles based on their affiliations and only keep the one with the highest score. Overall, we designed our scheme to minimize false matchings between sources rather than maximizing the number of profiles that would be matched; manual inspection shows that it is very robust to homonymy problems in the matching between the two sources. The proportion of individuals that could be matched with a digital profile for subgroups can be found in Appendix B.
For each profile linked to an individual, we obtain a list of dated professional positions that are shared between the public and private sectors (or “spheres”, see Daho and Gally (2018); France and Vauchez (2017)). We classify organization names between public and private positions, using a rule-based method, progressively adding filters to capture all public organizations. This strategy enabled us to cover manually the entire sample. Companies entirely owned by the state (e.g. La Poste, France Television) are also coded as public, alongside companies with the “EPIC” status (e.g. SNCF, OSEO). Contrary to Daho and Gally (2018), companies that only have the state as a minority shareholder (e.g. Renault, TotalEnergies) are coded as private. In addition, positions at Universities and research bodies are not coded as either public or private: it appears they correspond to associate teaching positions that do not constitute the main occupation for people in our population, and so we leave them out. This reasoning leads us to also not code implication in a political campaign, as this is often superposed to other engagements. The same goes for positions in associations, naturally including charities, but also positions in unions, which include corporate representation bodies such as Medef. Finally, to ensure homogeneity with Journal Officiel, we do not include positions in world organizations (e.g. World Health Organization). This strategy of not coding uncertain cases is enabled by our modeling approach, as it doesn’t require full trajectories, and reduces the impact of such decisions on results. Without claiming this coding scheme is perfect, these decisions only have a minor impact on the substantial research findings, and all 1488 coding filters will be made fully accessible in the appendix for re-use.
A.3 Exact perimeter for organizational belonging
Readers interested in the functioning of the French “Grands Corps” will probably be interested in our precise identification of traces that define membership in any of them. Our definition is built in an empirical manner even if there are two defining principles. On the one hand, the definition is heavily constrained by the nature of the archive and what it documents, and on the other hand, it is based on a continuous discussion with sociologists and political scientists so the criterion adopted here would match usual accounts, and avoid being overly broad or narrow.
Concretely, we begin by surveying the Steinertriples API for the different groups we are interested in. We first filter out traces depending on the organization and their nature, using a rule-based approach summarized in the table (0 meaning we filter out). This was defined by discussing with field experts with concrete examples of the different cases.
Table 4: Type of traces used to define membership by organization
short_name affectation nomination intégration titularisation
cab 0 1 0 0
ccomptes 1 1 1 0
cdiplo 0 1 0 0
ce 1 1 0 1
cprefet 0 1 0 1
csprefet 0 1 0 1
dgbudget 1 1 1 1
dgfip 1 1 1 1
dgtresor 1 1 1 1
ena 0 1 0 0
iga 1 1 1 1
igas 1 1 1 1
igf 1 1 1 1
insee 0 1 0 0
After this first filter, we apply specialized regex filters to precisely filter the population so it resembles as much as possible expert interpretations.
  • For ccomptes, traces that are not affectation, we only keep matches of (auditeur)|(auditrice)|(conseiller)|(conseillère)|(président)|(magistrat). We include the array of elite civil servants working at the Court.
  • For insee, we only keep traces that match (administrat)|(inspect)|(direct) and do not match (regiss)|(élève). We thus focus on members of the legal insee grand corps.
  • For ce, traces that are not affectation, we only keep traces matching (auditeur)| (auditrice)|(conseiller)|(conseillère)|(maître). We focus on Conseillers d’États and auditors.
  • igas, iga, igf, for nomination traces, we only keep matches of (inspecteur) |(inspectrice). We thus focus on inspectors.
  • For dgtresor and dgfip, nomination traces, we only keep matches of (directeur) |(directrice)|(chef)|(administrat)|(délégué)|(attaché)|(conseiller). We also remove nomination traces matching (commission)|(expert)|(régisseur)| (suppléant). For dgfip, we also remove traces that document being a member of a cabinet. We thus mostly focus on directors and sub-directors of services and economics experts.
  • For ENA, we only keep traces of admitted students.
After applying this filter, we then remove nomination traces matching (membre) |(représentant) which correspond to internal elections.
Appendix B Descriptive measures
Table 5: Number of students by year of admission.
Year Count Year Count Year Count
1990 89 2000 107 2010 76
1991 97 2001 112 2011 74
1992 81 2002 135 2012 80
1993 97 2003 112 2013 77
1994 104 2004 99 2014 75
1995 95 2005 90 2015 79
1996 96 2006 83 2016 87
1997 96 2007 87 2017 85
1998 95 2008 78 2018 77
1999 98 2009 79 2019 75
Table 6: Number of individuals affiliated by service (1990-2022).
group count group count
insee 1164 igf 455
cdiplo 914 igas 402
cprefet 838 dgfip 373
ccomptes 629 iga 203
ce 626 dgtresor 171
Table 7: Proportion of individuals by group with a matched profile.
group prop group prop
ena 55.2 ce 37.9
dgtresor 54.4 insee 33.2
igf 52.3 dgfip 32.4
igas 44.5 cdiplo 26.8
ccomptes 41.5 cprefet 24.7
iga 37.9
Table 8: Proportion of men and ENA graduates by organization.
group prop_men prop_ena group prop_men prop_ena
cprefet 87.6 11.7 dgfip 72.4 11.8
cdiplo 82.7 8.4 iga 71.4 31.0
dgtresor 80.7 26.3 ce 66.3 34.5
ccomptes 78.2 31.3 insee 64.8 0.3
igf 75.8 37.8 igas 61.5 29.8
Appendix C Additional robustness checks
C.1 Choice of $\varphi$
In our implementation, we include the covariates $A(1)$ and $A(\varphi)$ with $\varphi=0.8$. We performed additional tests with different weighting parameters $\varphi$ for the main model of interest (study 1, model 3) $\varphi=0.6,0.7,0.8,0.9$, and compared with the reconstruction loss criterion ($0.222,0.239,0.216,0.229$). The best-performing model , is when choosing $\varphi=0.8$.
We note that the inference behaves in a very similar manner when setting $\varphi=0.6,0.7,0.8$, but changes when setting $\varphi=0.9$. When setting $\varphi=0.9$, $A(\varphi)$ and $A(1)$ are very correlated (correlation $0.93$) and redundant, and the model is not able to pick up on medium-term behavior of the trace-emission process, which was the reason why we introduced $A(\varphi),\varphi<1$ in the first place, with some notable change in the parameter posteriors. Based on this, and since other variables already have the purpose of tracking small-term and long-term changes, we argue that one should avoid setting $\varphi$ too high (or too low).
For additional verification, we also constructed a model that includes multiple values of $\varphi$ ($\varphi=0.6,0.8$), in addition to $\varphi=1$ (it is not possible to include all $\varphi=0.6$ to $0.9$ because it leads to an ill-conditioned linear system). The reconstruction loss is very close to the original model ($0.216$), and the parameters inferred are also very similar, with no change in substantive conclusions.
C.2 Data heterogeneity concerns and modular inference
In our inference procedure, we use both values of $(X_{t}^{i})$ which have been observed (on LinkedIn), and values which have been simulated from the appropriate conditional; both types of values will be used to infer the parameters of interest. As our two data sources are of very different nature, the heterogeneity it may create can be disquieting. Even in the slightest misspecification case, simulated data could be a lot less trustworthy than observed data, and errors on the former could be propagated to the entire estimation procedure. In addition, even if we attempted to classify digital data so it would match as best as possible administrative accounts, some residual artefactual differences could still be observed.
This problem of having multiple, heterogeneous data modules is common and has been discussed in the Bayesian literature Jacob et al. (2017). For each configuration, we consider two settings: (a) all values are treated equally; (b) simulated values are given a weight of 0.01, whereas observed values are given a weight of 1. This second strategy effectively grants all importance to individuals who have a digital profile.
Simulations in modular inference control condition leads to a similar conclusion on $\gamma^{1}$, with a posterior on $\gamma^{0}$ a little more shifted to the negative side.
Appendix D Digital appendix and data availibility
The scripts and original data are available online at https://github.com/theovoldoire/Hidden_Careers_Public. The MCMC output for all models mentioned in the paper is available online at https://zenodo.org/doi/10.5281/zenodo.10365276.